# Embedding Similarity Guided License Plate Super Resolution

### v2

@@ Abstract
Super-resolution (SR) techniques play a pivotal role in enhancing the quality of low-resolution images, particularly for applications such as security and surveillance, where accurate license plate recognition is crucial. This study proposes a novel framework that combines pixel-based loss with embedding similarity learning to address the unique challenges of license plate super-resolution (LPSR). The introduced pixel and embedding consistency loss (PECL) integrates a Siamese network and applies contrastive loss to force embedding similarities to improve perceptual and structural fidelity. By effectively balancing pixel-wise accuracy with embedding-level consistency, the framework achieves superior alignment of fine-grained features between high-resolution (HR) and super-resolved (SR) license plates. Extensive experiments on the CCPD dataset validate the efficacy of the proposed framework, demonstrating consistent improvements over state-of-the-art methods in terms of $PSNR_{RGB}$, $PSNR_{Y}$ and optical character recognition (OCR) accuracy. These results highlight the potential of embedding similarity learning to advance both perceptual quality and task-specific performance in extreme super-resolution scenarios.
@@ 1 Introduction
In this work, we address the limitations of existing LPSR methods by focusing on extreme super-resolution scenarios with a challenging scaling factor of x8. To tackle the reconstruction of severely degraded license plates with minimal pixel information, we propose a deep learning framework that combines residual dense blocks (RDBs) and channel attention mechanisms to enhance visual quality and textual detail recovery. Our approach integrates pixel-level and embedding-level losses, implemented through a Siamese network to align embeddings of high-resolution (HR) and super-resolution (SR) images. By doing so, the character recovery is improved along with the overall fidelity. Besides, the training strategy based on embedding similarity utilizes contrastive loss [12] to minimize the discrepancy between HR and SR embeddings. This strategy preserves fine details critical for character recognition. Extensive experiments conducted on the CCPD dataset [47], which includes real-world license plate images under diverse conditions, demonstrate the effectiveness and robustness of the proposed method. The primary contributions of this work can be summarized as:
@@
    We developed a deep learning framework for extreme LPSR with a scaling factor of x8, leveraging residual dense blocks and channel attention mechanisms to enhance visual quality and recover fine details.
@@
    We conducted a comprehensive evaluation of the proposed method on the CCPD dataset, demonstrating its robustness and effectiveness across diverse real-world conditions.
@@ 2.1 Single image super-resolution
Despite these advancements, standard SISR models often struggle when applied to domain-specific tasks such as license plate super-resolution (LPSR), where the primary goal is not just to improve image fidelity, but also to preserve critical textual and structural information that is vital for recognition tasks. Indeed, the challenges are more domain-specific and tied to the unique visual characteristics of license plate images, such as small fonts, varying lighting conditions, and complex backgrounds. Traditional SISR models, when directly applied to license plates, tend to fail in recovering the fine-grained details required for character recognition, particularly when dealing with extreme scaling factors such as x8 or higher.
@@ 2.2 License plate super-resolution
Despite these advances, several challenges remain in LPSR research. First, many existing approaches are designed for moderate upscaling factors (i.e. x2 or x4) and fail to generalize to extreme cases such as x8 and beyond, where the license plate details are severely degraded. Moreover, current LPSR models often struggle to balance the trade-off between improving perceptual quality and preserving critical textual information. While OCR-guided loss functions have shown promise in mitigating this issue, there is still room for improvement, particularly in cases where the input LR images suffer from extreme distortions, such as motion blur or severe compression artifacts.
@@ 3.1 Problem formulation
In our method, we model the problem using a patch-based approach. Rather than processing entire HR and LR images, we operate on smaller patches extracted from these images. Formally, let $\mathbf{I}^{HR}\in\mathbb{R}^{H\times W\times C}$ and $\mathbf{I}^{LR}\in\mathbb{R}^{h\times w\times C}$ represent the HR and LR images respectively, where $H>h$ and $W>w$. Instead of directly mapping the whole LR image to its HR counterpart, we extract small overlapping patches $\mathbf{P}^{HR}_{i}\in\mathbb{R}^{p\times p\times C}$ and $\mathbf{P}^{LR}_{i}\in\mathbb{R}^{p^{\prime}\times p^{\prime}\times C}$ from both $\mathbf{I}^{HR}$ and $\mathcal{I}^{LR}$.
@@ 3.2 LPSR model overview
To achieve high-quality super-resolution, the proposed model is built upon foundational observations derived from state-of-the-art Single Image Super Resolution (SISR) techniques, specifically incorporating Residual Dense Blocks (RDBs) [51] and channel attention mechanisms [4]. The architecture is designed to progressively refine low-resolution inputs $\mathbf{P}_{\text{LR}}\in\mathbb{R}^{3\times W\times H}$ into high-resolution outputs $\mathbf{P}_{\text{SR}}\in\mathbb{R}^{3\times W\times H}$, where $W$ and $H$ denote the width and height, respectively, through a series of strategically implemented processing stages.
@@ 4.1 Datasets and Implementation Details
Dataset: We use the Chinese city parking dataset (CCPD) [47], which contains over 200k images of license plates captured in various real-world conditions, including different angles, distances, and lighting scenarios. This dataset is particularly suitable for our study as it provides diverse examples of license plates that challenge traditional recognition systems. We select 100k from the CCPD dataset to train the model and 1k for testing and validation. This selection ensures the model has sufficient data for learning, while maintaining a separate subset for unbiased performance evaluation. Each image is segmented into patches of size $64\times 64$ pixels.
The training dataset exhibits diverse capturing conditions, as illustrated in Fig. 4, including variations in angles, distances, and lighting. Additionally, many of the HR images suffer from significant distortions, further complicating the super-resolution task and emphasizing the need for robust models capable of handling degraded inputs.
@@
As described in Sec.3.2, the LPSR model processes input patches instead of full images. This patch-based design is a widely adopted practice in super-resolution (SR) tasks[6, 27, 45], as it enhances computational efficiency while enabling the model to focus on localized features. Moreover, it effectively handles the inherent variability in resolution and aspect ratios of real-world license plate images, ensuring robust generalization across diverse input conditions. Fig. 5 showcases sample pairs of high-resolution (HR) and low-resolution (LR) patches. Given that the original images already exhibit visual degradations, the degradation process $\mathcal{D}$ is limited to downscaling. In this work, we address the challenging task of x8 scaling, categorized as extreme super-resolution due to the significant loss of visual information in LR inputs.
@@
Figure 5: Examples of training patches and the degradation process applied, highlighting the downscaling operation used to simulate low-resolution inputs.
@@
Implementation Details: The LPSR model is implemented using the PyTorch library [2] and trained on a server equipped with an Intel Xeon Silver 4208 2.1GHz CPU, 192GB of RAM, and an Nvidia Tesla V100S GPU with 32GB of memory. We train the model for 1000k iterations, a batch size of 128. We employ the Adam optimizer [19] to update the parameters of the model, with an initial learning rate $41e-$ 4. To adaptively adjust the learning rate throughout the training process, we utilize a cosine annealing learning rate scheduler [30]. The scheduler gradually decreases the learning rate over the iterations following a cosine curve. This dynamic adjustment aids in fine-tuning the model’s parameters as training progresses.
@@
Evaluation criteria: The evaluation of LPSR models focuses on two key aspects: visual quality and character recognition performance. Visual quality assesses how closely the SR images resemble the ground truth HR images, ensuring the restored images are perceptually faithful to the originals. For this purpose, peak signal-to-noise ratio (PSNR) is measured on both the full-color RGB channels ($PSNR_{RGB}$) and the luminance channel ($PSNR_{Y}$), with the latter being particularly significant due to its strong correlation with human visual perception. In addition to visual fidelity, the effectiveness of the super-resolution model in enhancing character readability is evaluated through OCR. PaddleOCR [43], a widely-used open-source OCR system [40, 41, 32, 21], is employed to read characters on the SR images. The performance is quantified using the confidence scores provided by PaddleOCR, reflecting the accuracy and reliability of character recognition. This dual evaluation framework ensures a rebust assessment of the efficiency of the LPSR model.
@@ 4.2.1 Performance comparison
[…]
Bicubic - 18.53 ($\pm$ 2.78) 18.70 ($\pm$ 2.82) 0.697 ($\pm$ 0.19)
SRCNN [6] 57K 19.57 ($\pm$ 2.72) 19.84 ($\pm$ 2.78) 0.665 ($\pm$ 0.21)
MSRN [25] 6M 19.78 ($\pm$ 2.74) 19.84 ($\pm$ 2.82) 0.656 ($\pm$ 0.20)
ESPCN [42] 800K 23.52 ($\pm$ 2.47) 24.11 ($\pm$ 0.61) 0.848 ($\pm$ 0.14)
ESRGAN [46] 16M 18.76 ($\pm$ 2.26) 19.01 ($\pm$ 2.33) 0.812 ($\pm$ 0.17)
TBSRN [5] 12M 23.76 ($\pm$ 2.48) 24.42 ($\pm$ 2.63) 0.855 ($\pm$ 0.15)
SwinIR [27] 11M 23.56 ($\pm$ 2.50) 24.18 ($\pm$ 2.65) 0.850 ($\pm$ 0.14)
Ours 2M 25.13 ($\pm$ 2.46) 25.92 ($\pm$ 2.62) 0.852 ($\pm$ 0.13)
[…]
Traditional methods, such as Bicubic, SRCNN, and MSRN, exhibit the weakest performance across all metrics. Bicubic interpolation’s simplicity leads to significant oversmoothing, which degrades both image fidelity and text clarity. SRCNN and MSRN, though early milestones in deep-learning-based SR, lack advanced features like attention mechanisms or perceptual losses, making them less effective for domain-specific tasks like license plate recognition.
@@
Figure 6: Performance in terms of PSNR vs. the number of iterations (on a logarithmic scale) required for convergence during training.
The superiority of the proposed model stems from its ability to dynamically balance pixel-level fidelity and embedding similarity, effectively addressing the trade-off between visual quality and text preservation. Its design demonstrates the importance of combining domain-specific considerations with state-of-the-art architectural and loss functions to achieve robust performance in challenging SR scenarios.
@@
We further analyzed the performances in terms of PSNR on the testing set versus the number of iterations required for each model to converge. All models were trained for 1000k iterations, with the best version saved at the highest PSNR during training. The plot in Fig. 6 illustrates the PSNR vs. number of iterations (on a logarithmic scale). A clear correlation between the number of iterations and the resulting PSNR values can be observed, except for SRCNN.
@@
Figure 7: Qualitative comparison with state-of-the-art methods on various samples, taken under different conditions.
@@
To further compare the performance of the proposed model with state-of-the-art methods, we provide a qualitative comparison in Fig. 7. This comparison showcases the visual differences in SR outputs across various models compared to the proposed one, including Bicubic interpolation, SRCNN, MSRN, ESPCN, ESRGAN, TBSRN, and SwinIR. The results indicate significant variability in the restoration of fine details and overall visual fidelity.
Bicubic interpolation serves as a baseline, delivering overly smooth images with a lack of texture recovery, underscoring its inability to reconstruct high-frequency details. SRCNN, being one of the earliest SR models, provides slightly sharper results but still struggles with generating realistic textures, leading to noticeable artifacts in regions with fine details. MSRN and ESPCN demonstrate improved performance, with MSRN showing better preservation of structural information and ESPCN excelling in edge sharpness. However, both models exhibit some limitations in producing natural textures.
@@
ESRGAN’s outputs, known for its GAN-based approach, introduce more realistic details and textures but often at the cost of over-enhanced or unnatural artifacts in certain areas, especially on less complex regions. TBSRN and SwinIR deliver a strong balance between detail recovery and artifact suppression, with SwinIR slightly outperforming TBSRN in preserving structural consistency across diverse regions of the images. The proposed model achieves the best results overall, demonstrating superior texture restoration, edge sharpness, and fidelity to high-resolution ground truths, with minimal artifacts and a more natural appearance.

### v3

@@ Abstract
Super-resolution (SR) techniques play a pivotal role in enhancing the quality of low-resolution images, particularly for applications such as security and surveillance, where accurate license plate recognition is crucial. This study proposes a novel framework that combines pixel-based loss with embedding similarity learning to address the unique challenges of license plate super-resolution (LPSR). The introduced pixel and embedding consistency loss (PECL) integrates a Siamese network and applies contrastive loss to force embedding similarities to improve perceptual and structural fidelity. By effectively balancing pixel-wise accuracy with embedding-level consistency, the framework achieves superior alignment of fine-grained features between high-resolution (HR) and super-resolved (SR) license plates. Extensive experiments on the CCPD and PKU dataset validate the efficacy of the proposed framework, demonstrating consistent improvements over state-of-the-art methods in terms of PSNR, SSIM, LPIPS, and optical character recognition (OCR) accuracy. These results highlight the potential of embedding similarity learning to advance both perceptual quality and task-specific performance in extreme super-resolution scenarios.
@@ 1 Introduction
In this work, we address the limitations of existing LPSR methods by focusing on extreme super-resolution scenarios with a challenging scaling factor of $\times 8$. To tackle the reconstruction of severely degraded license plates with minimal pixel information, we propose a deep learning framework that combines residual dense blocks (RDBs) and channel attention mechanisms to enhance visual quality and textual detail recovery. Our approach integrates pixel-level and embedding-level losses, implemented using a Siamese network to align embeddings of high-resolution (HR) and super-resolution (SR) images. By doing so, the character recovery is improved along with the overall fidelity. Besides, the training strategy based on embedding similarity utilizes contrastive loss [13] to minimize the discrepancy between HR and SR embeddings. This strategy preserves fine details that are critical for character recognition. Extensive experiments conducted on both CCPD [51] and PKU [54] datasets, containing real-world license plate images under diverse conditions, demonstrate the effectiveness and robustness of the proposed method. The primary contributions of this work can be summarized as:
@@
    We developed a deep learning framework for extreme LPSR with a scaling factor of $\times 8$, leveraging residual dense blocks and channel attention mechanisms to enhance visual quality and recover fine details.
@@
    We conducted a comprehensive evaluation of the proposed method on both CCPD and PKU datasets, demonstrating its robustness and effectiveness across diverse real-world conditions.
@@ 2.1 Single image super-resolution
Despite these advancements, standard SISR models often struggle when applied to domain-specific tasks such as license plate super-resolution (LPSR), where the primary goal is not just to improve image fidelity, but also to preserve critical textual and structural information that is vital for recognition tasks. Indeed, the challenges are more domain-specific and tied to the unique visual characteristics of license plate images, such as small fonts, varying lighting conditions, and complex backgrounds. Traditional SISR models, when directly applied to license plates, tend to fail in recovering the fine-grained details required for character recognition, particularly when dealing with extreme scaling factors such as $\times 8$ or higher.
@@ 2.2 License plate super-resolution
Despite these advances, several challenges remain in LPSR research. First, many existing approaches are designed for moderate upscaling factors (i.e. x2 or x4) and fail to generalize to extreme cases such as $\times 8$ and beyond, where the license plate details are severely degraded. Moreover, current LPSR models often struggle to balance the trade-off between improving perceptual quality and preserving critical textual information. While OCR-guided loss functions have shown promise in mitigating this issue, there is still room for improvement, particularly in cases where the input LR images suffer from extreme distortions, such as motion blur or severe compression artifacts.
@@ 3.1 Problem formulation
In our method, we model the problem using a patch-based approach. Rather than processing entire HR and LR images, we operate on smaller patches extracted from these images. Formally, let $\mathbf{I}^{HR}\in\mathbb{R}^{C\times H\times W}$ and $\mathbf{I}^{LR}\in\mathbb{R}^{C\times h\times w}$ represent the HR and LR images respectively, where $H>h$ and $W>w$. Instead of directly mapping the whole LR image to its HR counterpart, we extract small overlapping patches $\mathbf{P}^{HR}_{i}\in\mathbb{R}^{C\times p\times p}$ and $\mathbf{P}^{LR}_{i}\in\mathbb{R}^{C\times p^{\prime}\times p^{\prime}}$ from both $\mathbf{I}^{HR}$ and $\mathbf{I}^{LR}$.
@@ 3.2 LPSR model overview
To achieve high-quality super-resolution, the proposed model is built upon foundational observations derived from state-of-the-art Single Image Super Resolution (SISR) techniques, specifically incorporating residual dense blocks (RDBs) [57] and channel attention mechanisms [4]. The architecture is designed to progressively refine low-resolution inputs $\mathbf{P}_{\text{LR}}\in\mathbb{R}^{C\times W\times H}$ into high-resolution outputs $\mathbf{P}_{\text{SR}}\in\mathbb{R}^{C\times W\times H}$, where $C$, $W$, and $H$ denote the number of channels, width and height, respectively, through a series of strategically implemented processing stages.
@@ 4.1 Datasets and Implementation Details
Dataset: To assess the performance of our proposed approach across diverse real-world scenarios, we use two publicly available datasets: the Chinese City Parking Dataset (CCPD)[51] and the PKU Vehicle Dataset[54].
CCPD [51]:
This dataset comprises over 200k images of license plates captured under varying real-world conditions, including different angles, distances, and lighting scenarios. Its diversity makes it particularly well-suited for our study, as it presents significant challenges for traditional recognition systems. We use 100,000 images for training and 1k for testing and validation, ensuring well-balanced sets for both learning and evaluation.
PKU [54]:
This dataset consists of over 4k license plate images taken in diverse environments, such as highways under normal daylight and nightlight conditions and intersections with crosswalks during both day and night. We randomly select 3,5k images for training and allocate the remaining for testing, ensuring representative data distributions.
The training datasets exhibit diverse capturing conditions, as illustrated in Fig. 4, including variations in angles, distances, and lighting. Additionally, many of the HR images suffer from significant distortions, further complicating the super-resolution task and emphasizing the need for robust models capable of handling degraded inputs.
@@
As described in Sec.3.2, the LPSR model processes input patches instead of full images. This patch-based design is a widely adopted practice in super-resolution (SR) tasks[7, 28, 47], as it enhances computational efficiency while enabling the model to focus on localized features. Moreover, it effectively handles the inherent variability in resolution and aspect ratios of real-world license plate images, ensuring robust generalization across diverse input conditions. Fig. 5 showcases sample pairs of high-resolution (HR) and low-resolution (LR) patches. Given that the original images already exhibit visual degradations, the degradation process $\mathcal{D}$ is limited to downscaling. In this work, we address the challenging task of $\times 8$ scaling, categorized as extreme super-resolution due to the significant loss of visual information in LR inputs. For this study, the patch size is set $64\times 64$ pixels.
@@
Figure 5: Examples of training patches and the applied degradation process, highlighting the downscaling operation used to simulate low-resolution inputs. (left) example from CCPD and (right) from PKU.
@@
Implementation Details: The LPSR model is implemented using the PyTorch library [2] and trained on a server equipped with an Intel Xeon Silver 4208 2.1GHz CPU, 192GB of RAM, and an Nvidia Tesla V100S GPU with 32GB of memory. We train the model for 1000k iterations with a batch size of 128. We employ the Adam optimizer [20] to update the parameters of the model, with an initial learning rate 1e41e-4. To adaptively adjust the learning rate throughout the training process, we utilize a cosine annealing learning rate scheduler [31]. The scheduler gradually decreases the learning rate over the iterations following a cosine curve. This dynamic adjustment aids in fine-tuning the model’s parameters as training progresses.
@@
Evaluation criteria: The evaluation of LPSR models focuses on two key aspects: visual quality and character recognition performance. Visual quality assesses how closely the SR images resemble the ground truth HR images, ensuring the restored images are perceptually faithful to the originals. For this purpose, we use peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) [49], and learned perceptual image patch similarity (LPIPS) [55]. PSNR and SSIM are measured on both full-color RGB channels (denoted as $PSNR$ and $SSIM$) and luminance channel (denoted as $PSNR_{Y}$ and $SSIM_{Y}$), with the latter being particularly significant due to its strong correlation with human visual perception. In addition to visual fidelity, the effectiveness of the super-resolution model in enhancing character readability is evaluated using an OCR. PaddleOCR [44], a widely-used open-source OCR system [41, 42, 33, 22], is employed to read characters on the SR images. The performance is quantified using several metrics, including 1) exact match accuracy (EMA) to quantify the percentage of cases where the predicted LP text exactly matches the ground truth, 2) The Average Levenshtein similarity (L-sim) [46], which provides a measure of partial correctness by computing the edit distance between predictions and references, and 3) character error rate (CER) and word error rate (WER) assess the frequency of misrecognized characters and words, respectively. In addition, to capture performance at a more granular level, average character-level precision (Prec.), recall, and F1-scores (F1-S) are computed, reflecting the accuracy and consistency of individual character predictions. As these metrics depend on ground truth annotations, the OCR-based evaluation is conducted exclusively on the CCPD dataset as it provides such annotations.
@@ 4.2.1 Performance comparison
The performance of the proposed SR model is evaluated in comparison to the baseline Bicubic and SOTA methods, including SRCNN [7], MSRN [26], ESPCN [43], ESRGAN [48], TBSRN [5], and SwinIR [28]. Table 2 presents the results of this comparison, where the median values, along with the standard deviations, are reported over the testing set. In addition to quality scores, Table 2 reports the complexity analysis quantified by means of number of parameters (#P) and floating point operations per second (GFLOPs).
Table 1: Performance comparison of the proposed model with various SR methods on the CCPD dataset [51]. The median ($\pm$ standard deviation) over the testing set is reported. The best and second-best performances are respectively highlighted in bold red and bold blue.
SR Method PSNR/PSNRy ($\uparrow$) SSIM/SSIMy ($\uparrow$) LPIPS ($\downarrow$) #P / GFLOPs
Bicubic 18.53 ($\pm$ 2.78) / 18.70 ($\pm$ 2.82) 0.4504 ($\pm$ 0.11) / 0.4619 ($\pm$ 0.12) 0.395 ($\pm$ 0.07) -
SRCNN [7] 19.57 ($\pm$ 2.72) / 19.84 ($\pm$ 2.78) 0.5426 ($\pm$ 0.10) / 0.5675 ($\pm$ 0.11) 0.319 ($\pm$ 0.11) 57K / 0.28
MSRN [26] 19.78 ($\pm$ 2.74) / 19.84 ($\pm$ 2.82) 0.5069 ($\pm$ 0.10) / 0.5237 ($\pm$ 0.11) 0.362 ($\pm$ 0.07) 6M / 15.11
ESPCN [43] 23.52 ($\pm$ 2.47) / 24.11 ($\pm$ 0.61) 0.7383 ($\pm$ 0.06) / 0.7750 ($\pm$ 0.07) 0.146 ($\pm$ 0.04) 800K / 14.50
ESRGAN [48] 18.76 ($\pm$ 2.26) / 19.01 ($\pm$ 2.33) 0.5709 ($\pm$ 0.08) / 0.5990 ($\pm$ 0.08) 0.122 ($\pm$ 0.04) 16M / 17.64
TBSRN [5] 23.76 ($\pm$ 2.48) / 24.42 ($\pm$ 2.63) 0.7480 ($\pm$ 0.07) / 0.7854 ($\pm$ 0.08) 0.143 ($\pm$ 0.05) 12M / 18.49
SwinIR [28] 23.56 ($\pm$ 2.50) / 24.18 ($\pm$ 2.65) 0.7477 ($\pm$ 0.06) / 0.7816 ($\pm$ 0.07) 0.147 ($\pm$ 0.05) 11M / 46.45
Ours 25.13 ($\pm$ 2.46) / 25.92 ($\pm$ 2.62) 0.8127 ($\pm$ 0.06) / 0.8458 ($\pm$ 0.06) 0.106 ($\pm$ 0.04) 1.9M / 13.35
Table 2: Performance comparison of the proposed model with various SR methods on the PKU dataset [54]. The median ($\pm$ standard deviation) over the testing set is reported. The best and second-best performances are respectively highlighted in bold red and bold blue.
SR Method PSNR/PSNRy ($\uparrow$) SSIM/SSIMy ($\uparrow$) LPIPS ($\downarrow$) #P / GFLOPs
Bicubic 15.89 ($\pm$ 1.95) / 15.98 ($\pm$ 1.98) 0.2832 ($\pm$ 0.09) / 0.2839 ($\pm$ 0.09) 0.578 ($\pm$ 0.07) -
SRCNN [7] 15.37 ($\pm$ 1.95) / 15.41 ($\pm$ 1.92) 0.3267 ($\pm$ 0.09) / 0.3341 ($\pm$ 0.10) 0.465 ($\pm$ 0.08) 57K / 0.28
MSRN [26] 15.58 ($\pm$ 2.17) / 15.61 ($\pm$ 2.15) 0.3132 ($\pm$ 0.10) / 0.3185 ($\pm$ 0.11) 0.4442 ($\pm$ 0.09) 6M / 15.11
ESPCN [43] 17.52 ($\pm$ 2.99) / 17.61 ($\pm$ 2.15) 0.5492 ($\pm$ 0.12) / 0.5743 ($\pm$ 0.12) 0.199 ($\pm$ 0.06) 800K / 14.50
ESRGAN [48] 18.74 ($\pm$ 3.59) / 18.92 ($\pm$ 3.69) 0.6796 ($\pm$ 0.13) / 0.7014 ($\pm$ 0.13) 0.110 ($\pm$ 0.05) 16M / 17.64
TBSRN [5] 17.31 ($\pm$ 3.17) / 17.47 ($\pm$ 3.22) 0.5578 ($\pm$ 0.13) / 0.5813 ($\pm$ 0.13) 0.174 ($\pm$ 0.06) 12M / 18.49
SwinIR [28] 17.39 ($\pm$ 3.26) / 17.62 ($\pm$ 3.31) 0.5414 ($\pm$ 0.14) / 0.5672 ($\pm$ 0.15) 0.188 ($\pm$ 0.06) 11M / 46.45
Ours 19.26 ($\pm$3.35) / 19.46 ($\pm$3.77) 0.7009 ($\pm$0.12) / 0.7222 ($\pm$0.12) 0.101 ($\pm$0.04) 1.9M / 13.35
Figure 6: GFLOPs vs. PSNR on the CCPD dataset [51] for the proposed method compared to SOTA models. The scatter plot visualizes the trade-off between computational complexity (measured in GFLOPs) and reconstruction quality (measured in PSNR) for different SR models. Each point represents a model, with a radius proportional to the square root of its GFLOPs. The GFLOPs were computed using $64\times 64$ input images and averaged over 10 iterations for consistency.
Quantitative comparison: On the CCPD dataset, the performance trend clearly favors advanced deep learning methods over traditional interpolation and early CNN-based approaches. Bicubic interpolation, along with SRCNN and MSRN, delivers modest PSNR values in the range of 18–20 dB and relatively low SSIM scores, reflecting their limited capacity to reconstruct fine details and textures from low-resolution inputs. In contrast, methods such as ESPCN, ESRGAN, TBSRN, and SwinIR exhibit marked improvements, with ESPCN achieving a PSNR of 23.52 dB and TBSRN and SwinIR reaching similar high-quality scores. Notably, the proposed model surpasses all competitors by achieving a PSNR of 25.13 dB (with a PSNRy of 25.92 dB) and an SSIM of 0.8127/0.8458, along with the lowest LPIPS of 0.106. These superior scores, combined with a lightweight architecture (1.9M parameters and 13.35 GFLOPs), underscore the model’s ability to enhance perceptual quality, preserve structural details, and effectively handle the diverse imaging conditions inherent in the large-scale CCPD dataset.
In comparison, the PKU dataset, characterized by its smaller scale and generally lower inherent image quality, presents a more challenging testbed for super-resolution methods. On this dataset, methods such as Bicubic, SRCNN, and MSRN yield even lower PSNR values (around 15–16 dB) and SSIM scores, while modern deep learning techniques show improved performance. For instance, ESPCN reaches a PSNR of approximately 17.5 dB and an SSIM close to 0.55, and ESRGAN, highlighted as the second-best performance, attains a PSNR of 18.74 dB and an SSIM of 0.6796. However, the proposed model again takes the lead with a PSNR of 19.26 dB (PSNRy of 19.46 dB), an SSIM of 0.7009/0.7222, and the lowest LPIPS of 0.101. This consistency in outperforming state-of-the-art methods on both datasets demonstrates the robustness and adaptability of the proposed approach across varying scales and quality levels, effectively balancing high-fidelity reconstruction with computational efficiency.
Refer to caption Refer to caption
@@
Figure 7: Performance in terms of PSNR vs. the number of iterations (on a logarithmic scale) required for convergence during training. (left) CCPD dataset [51] and (right) PKU dataset [54].
Additionally, the analysis of computational complexity in terms of GFLOPs versus PSNR, as depicted in Fig. 6, reveals that higher computational complexity does not always guarantee better performance. While models like SwinIR (46.45 GFLOPs, 23.56 PSNR) exhibit high computational demands, they do not achieve the best PSNR. In contrast, TBSRN (18.49 GFLOPs, 23.76 PSNR) and ESPCN (14.5 GFLOPs, 23.52 PSNR) reach comparable if not better performance with significantly lower GFLOPs, highlighting the efficiency of certain architectures. Additionally, the trend suggests diminishing returns beyond approximately 15-20 GFLOPs, where increased computation does not necessarily yield substantial PSNR improvements. For example, ESRGAN (17.64 GFLOPs, 18.76 PSNR) has high computational cost but relatively low PSNR, likely due to its focus on perceptual quality rather than pixel-wise accuracy. Notably, our model (13.35 GFLOPs, 25.13 PSNR) achieves the highest PSNR while maintaining moderate computational complexity, indicating an efficient design. This suggests that optimized network architectures and better feature representations can significantly improve performance without excessive computational cost.
@@
We further analyzed the performances in terms of PSNR on the testing set versus the number of iterations required for each model to converge. All models were trained for 1000k iterations, with the best version saved at the highest PSNR during training. The plot in Fig. 7 illustrates the PSNR vs. number of iterations (on a logarithmic scale) on both dataset. A clear correlation between the number of iterations and the resulting PSNR values can be observed, except for SRCNN.
@@
Figure 8: Qualitative comparison with state-of-the-art methods on various samples from the CCPD dataset [51], taken under different conditions.
Figure 9: Qualitative comparison with state-of-the-art methods on various samples from the PKU dataset [54], taken under different conditions.
@@
Qualitative comparison: To further compare the performance of the proposed model with state-of-the-art methods, we provide a qualitative comparison in Fig. 8 for CCPD and Fig. 9 for PKU dataset. This comparison showcases the visual differences in SR outputs across various models compared to the proposed one, including Bicubic interpolation, SRCNN, MSRN, ESPCN, ESRGAN, TBSRN, and SwinIR. The results indicate significant variability in the restoration of fine details and overall visual fidelity.
@@
On the CCPD dataset, the Bicubic interpolation serves as a baseline, delivering overly smooth images with a lack of texture recovery, underscoring its inefficiency to reconstruct high-frequency details. SRCNN, being one of the earliest SR models, provides slightly sharper results but still struggles with generating realistic textures, leading to noticeable artifacts in regions with fine details. MSRN and ESPCN demonstrate improved performance, with MSRN showing better preservation of structural information and ESPCN excelling in edge sharpness. However, both models exhibit some limitations in producing natural textures. ESRGAN’s outputs, known for its GAN-based approach, introduce more realistic details and textures but often at the cost of over-enhanced or unnatural artifacts in certain areas, especially on less complex regions. TBSRN and SwinIR deliver a strong balance between detail recovery and artifact suppression, with SwinIR slightly outperforming TBSRN in preserving structural consistency across diverse regions of the images. The proposed model achieves the best results overall, demonstrating superior texture restoration, edge sharpness, and fidelity to high-resolution references, with minimal artifacts and a more natural appearance.
The qualitative comparison on the PKU dataset reveals the significant challenges posed by heavily degraded low-resolution inputs, where baseline methods such as Bicubic, SRCNN, and MSRN struggle to recover high-frequency details, resulting in outputs that are notably blurry and lack sharpness. ESPCN offers a modest improvement by slightly enhancing edge preservation. Methods with more advanced architectures using ViTs like TBSRN and SwinIR further reduce blurring and improve detail recovery. Notably, ESRGAN exhibits a marked improvement on the PKU dataset compared to its performance on the CCPD dataset, a difference that can be largely attributed to the inherent characteristics of PKU, i.e. its smaller scale and reduced variability make it more suitable to ESRGAN’s GAN-based enhancement of realistic textures. Despite this, the proposed model consistently stands out by generating images with sharp edges and rich textures that closely match the high-resolution references, again demonstrating its robust ability to reconstruct perceptually meaningful details even from severely degraded inputs.
Recognition performance analysis: We conduct a license plate recognition analysis using several metrics as described in Sec.4.1 to further analyze the efficiency of the proposed method in recovering task-aware characteristics that facilitate better recognition on the CCPD dataset. Table 3 presents a comprehensive comparison of text recognition performance on super-resolved license plate images across several methods on the CCPD datset. Bicubic interpolation, serving as a baseline, achieves an EMA of 23.70% and an average $L_{\text{sim}}$ of 60.88%, with high error rates (CER of 39.12% and WER of 76.30%), indicating its limited capacity to restore fine textual details. Early CNN-based methods like SRCNN and MSRN perform even lower, with EMA of 17.50% and 18.30% respectively, and similarly high error metrics, underscoring their challenges in capturing and reconstructing complex text structures. In contrast, methods leveraging deeper architectures and more advanced perceptual losses show marked improvements: ESPCN reaches an EMA of 56.60% with a CER of 15.02% and WER of 43.40%, while ESRGAN, despite its GAN-based approach, records a moderate performance with 43.00% EMA. Notably, TBSRN and SwinIR further enhance performance, achieving EMA of 57.40% and 58.10%, accompanied by lower error rates and improved character-level precision, recall, and F1-scores. The proposed method outperforms all competitors by reaching the highest EMA of 62.80% and the best average $L_{\text{sim}}$ of 88.38%, along with the lowest CER (11.62%) and WER (37.20%). Its superior character-level metrics, with precision, recall, and F1-scores exceeding 82%. These performances underscore its effectiveness in enhancing textual legibility and recognition accuracy. These results collectively demonstrate that the proposed approach offers a significant advancement in recovering perceptually meaningful details from degraded license plate images, thereby enabling more accurate text recognition compared to both conventional and state-of-the-art super-resolution methods.
Table 3: Text recognition performance on super-resolved license plates from the CCPD dataset. The best and second-best results are highlighted in bold red and bold blue, respectively.
[…]
Bicubic 23.70% 60.88% 39.12% 76.30% 49.47% 49.72% 49.36%
SRCNN [7] 17.50% 52.47% 47.53% 82.50% 40.37% 40.53% 40.21%
MSRN [26] 18.30% 55.90% 44.10% 81.70% 42.80% 43.09% 42.70%
ESPCN [43] 56.60% 84.98% 15.02% 43.40% 78.16% 78.41% 78.11%
ESRGAN [48] 43.00% 76.00% 24.00% 57.00% 67.15% 67.52% 67.09%
TBSRN [5] 57.40% 85.67% 14.33% 42.60% 79.09% 79.22% 79.00%
SwinIR [28] 58.10% 86.10% 13.90% 41.90% 79.63% 79.86% 79.57%
Ours 62.80% 88.38% 11.62% 37.20% 82.07% 82.20% 82.01%